OpenStudy (anonymous):

HELP WILL GIVE MEDAL Find an equation for the nth term of the sequence. -4, -16, -64, -256, ...

OpenStudy (trojanpoem):

a = -4 common ratio = r = -16/-4 = 4 nth term = ar^n = -4 * (4)^n = -4^n+1 for first term n = 0, for second n = 1 and so on if you want to use 1 as first term -4^n

OpenStudy (anonymous):

so an=-4x4^n-1?

OpenStudy (trojanpoem):

For case of first term has n = 1

OpenStudy (anonymous):

im confused now

OpenStudy (anonymous):

ohh you are ssayin an=-4x4^n because n=1

OpenStudy (trojanpoem):

This is a geometric series, you multiply each time times a constant in our case 4. you can use this formula r = x_{n+1} / x_{n} x indicate the term a = first term r is the ration between 2 consecutive terms to find the term a_{n} x_{n}= ar^n-1

OpenStudy (anonymous):

oh so i was right the second time

OpenStudy (trojanpoem):

I told you that it depends on how you count zero is the first term or 1 is the first term.

OpenStudy (trojanpoem):

To me, I used to count from zero.

OpenStudy (anonymous):

I think it is 0

OpenStudy (trojanpoem):

See the formula: x_{n} = ar^(n-1) put n = 1 x_{1} = ar^0 x_{n} = ar^n put n = 0 x_{n} = ar^0

OpenStudy (anonymous):

Oh i definitley got the answer now i get it, it is an=-4x4^n-1 thanks!

OpenStudy (trojanpoem):

You're welcome.